OpenStudy (anonymous):

i need help with (-3√20 * 5)/6 and if u can show your work

OpenStudy (chaise):

\[\frac{-3\sqrt{20\times5}}{6}\] Is this right, or is square root just on the 20?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

its on 20 times 5

OpenStudy (chaise):

Cool. (-3sqrt(20x5))/6 = -3sqrt(100)/6 = -3(10)/6 = -30/6 = -5

OpenStudy (anonymous):